For this and the next 5 questions, please refer to the following Table. Data in millions of dollars. What is the net operating profit after taxes (NOPAT) for YEAR Income Statement Year 1 Year 2 Sales Operating costs excl depreciation Depreciation EBIT Less: Interest EBT Taxes (40%) Net income to common Common dividends s 1,200.0 $1,000.0 1,020.0 850.0 30.0 25.0 125.0 21.7 20.2 104.8 51.341.9 77,0 62,9 150.0 128.3 60.5 ASSETS İt to save and submit. Click Save AllAnsuers to save all
media%2F84d%2F84d5680a-5a7e-4e99-921a-65
ites here, select then *, and drag to the Favorites Bar folder. Or import from ano CLAIMS Accounts payable $ 108.0 90.0 51.5 60.0 201.5 150.0 351.5 50.0 225.0208.5 275.0 258.5 672.0 610.0 Notes payable 67.0 72.0 247.0 150.0 397.0 50.0 Accruals Total current liabilities Long-term bonds Total debt Common stock (50 million shares) Retained earnings Common equity Total claims O$150 million $90 million $210 million $192 million None of the above
Calculate the net operating working capital for Year 1 O $150 million O $90 million O $210 million O $192 million O None of the above QUESTION 3 Calculate the net operating working capital for Year 2 $150 million $90 million $210 million $192 million None of the above
Calculate the total operating capital for Year 1 O s492 million O $58 million O $460 million O None of the above QUESTION 5 Calculate the total operating capital for Year 2 O $492 million O $58 million O $460 million O None of the above QUESTION 6 What is the free cash flow (FCF) for Year 2? $492 million $58 milion S460 millian al
Answer #1
Q-2) NOWC for Year 1 192
Q-3) NOWC for Year 2 210
Q-4) TOC for Year 1 492
Q-5) TOc for Year 2 460
Q-6) FCF for Year 2 15.1
Net operating working capital: Year 1 Year 2
Cash and cash equivalents 12 10
Accounts receivable 180 150
Inventories 180 200
Total operating current assets 372 360
Accounts payable 108 90
Accruals 72 60
Total operating current liabilities 180 150
Net operating working capital 192 210
Net plant and equipment 300 250
Total operating capital 492 460
NI 62.9
Add: Depreciation 25.0
Add: Interest 20.2
Operating cash flow 108.1
Less: Net capital spending = 300-250+25 = 75.0
Less: Change in NWC = (360-150)-(372-180) = 18.0
FCF 15.1
Options given for FCF are wrong